# Gaze-based dual resolution deep imitation learning for high-precision dexterous robot manipulation

### v3

@@ I Introduction
Figure 1: The proposed method can efficiently calculate a precise policy with both global features from peripheral vision (0(a)) and detailed visual information for the needle and thread from foveated vision (0(b)).
@@ III-C Recovery action
Figure 5: Example of task failure caused by the human operator. The operator failed to thread the needle (4(b)), recovered from the failure (4(c)), retried, and finally succeeded (4(d)).
@@ III-D Model architecture
By contrast, the recognition of the three-dimensional state of the thread and the needle is important for the processing of the foveated image. Spatial softmax is inadequate for inferring such information because the extracted feature is a two-dimensional coordinate. Therefore, the foveated image is processed using a series of convolutional neural networks and max-pooling layers with a stride of two (Fig. 5(b), 5(c)).
@@ IV-A Task setup
(b) Assessment of foveated vision Without gaze (image size: $128\times 72$) 93.75 $*{}^{*}$
Without gaze (image size: $1280\times 720$) $*{}^{*}$ $*{}^{*}$
(c) Assessment of action separation Without action separation $*{}^{*}$ $*{}^{*}$
Action speed threshold = 0.00208 $*{}^{*}$ $*{}^{*}$
@@
Action speed threshold = 0.00985 87.50 $*{}^{*}$
(d) Assessment of visuomotor control separation Fast-action with foveated vision only $*{}^{*}$ $*{}^{*}$
Fast-action with foveated + peripheral vision $*{}^{*}$ $*{}^{*}$
@@
No recovery step predictor 93.75 $*{}^{*}$
@@
(g) Assessment of stereo vision Left image only 68.75 $*{}^{*}$
@@
TABLE I: Needle threading result. The proposed method was significantly better than the task results marked using $*{}^{*}$ (chi-square test, $P<0.05$).

### v4

@@ I Introduction
Figure 1: The proposed method can efficiently calculate a precise policy with both global features from peripheral vision (1(a)) and detailed visual information for the needle and thread from foveated vision (1(b)).
@@ III-C Recovery action
Figure 5: Example of task failure caused by the human operator. The operator failed to thread the needle (5(b)), recovered from the failure (5(c)), retried, and finally succeeded (5(d)).
@@ III-D Model architecture
By contrast, the recognition of the three-dimensional state of the thread and the needle is important for the processing of the foveated image. Spatial softmax is inadequate for inferring such information because the extracted feature is a two-dimensional coordinate. Therefore, the foveated image is processed using a series of convolutional neural networks and max-pooling layers with a stride of two (Fig. 6(b), 6(c)).
@@ IV-A Task setup
(b) Assessment of foveated vision Without gaze (image size: $128\times 72$) 93.75 12.50
Without gaze (image size: $1280\times 720$) 0.00 0.00
(c) Assessment of action separation Without action separation 43.75 0.00
Action speed threshold = 0.00208 25.00 12.50
@@
Action speed threshold = 0.00985 87.50 43.75
(d) Assessment of visuomotor control separation Fast-action with foveated vision only 0.00 0.00
Fast-action with foveated + peripheral vision 6.25 0.00
@@
No recovery step predictor 93.75 31.25
@@
(g) Assessment of stereo vision Left image only 68.75 12.50
@@
TABLE I: Needle threading result. The proposed method was significantly better than the task results marked using (chi-square test, $P<0.05$).